OpenStudy (anonymous):

convert the following into an exponential equation: log(81)^3-1/4 a. 3^81=1/4 b. 81^4=3 c. 3^1/4=81 d. 81^1/4=3 e. none

jimthompson5910 (jim_thompson5910):

\[\Large \log_{81}(3)=\frac{1}{4}\] \[\Large 81^{\frac{1}{4}}=3\] So the answer is choice D

jimthompson5910 (jim_thompson5910):

I'm using the idea that \[\LARGE \log_{b}(x)=y\] can be written as \[\LARGE b^y=x\]